OpenStudy (anonymous):

solve 4 - 7x + 10 = 15 - 6x - 8

OpenStudy (anonymous):

try this one and tell me your answer then i compare with mine

OpenStudy (anonymous):

did u try?

OpenStudy (anonymous):

yes got it thanks

OpenStudy (anonymous):

man odubo they are messing with you . this answer is not right

OpenStudy (anonymous):

wat did u get?

OpenStudy (anonymous):

well maybe i made a mistake. let me try again

OpenStudy (anonymous):

\[4-7x+10=15-6x-8\] \[14-7x=7-6x\] \[14-7=-6x+7x\] \[7=x\]

OpenStudy (anonymous):

that is what i got. of course i didn't "move" anything to the right or left. i added and subtracted from both sides of the equation

OpenStudy (anonymous):

really today it is caveat emptor, even though there are no "buyers"

OpenStudy (anonymous):

I'm wrong, x does = 7, sorry